MICROWAVE SPECTRUM AND STRUCTURE OF A POLAR DIMER OF N_{2}O

\underline{NICHOLAS R. WALKER}, School of Chemistry, University of Bristol, Bristol, BS8 1TS, U.K.; AN-DREA J. MINEI, STEWART E. NOVICK, Department of Chemistry, Wesleyan University, Middletown, Connecticut 06459, U.S.A.; and ANTHONY C. LEGON, School of Chemistry, University of Bristol, Bristol, BS8 1TS, U.K..

Cavity Fourier-Transform microwave spectroscopy has been used to characterise a gas phase, polar dimer of N_{2}O. The polar (N_{2}O)_{2} unit is generated by co-expansion of a gas sample containing a small percentage of N_{2}O in a helium backing gas. Transitions in the pure rotational spectra of (^{15}N_{2}O)_{2}, (^{14}N^{15}NO)(^{15}N_{2}O), (^{14}N_{2}O)(^{15}N_{2}O) and (^{14}N_{2}O)_{2} are reported. The measured transitions of (^{15}N_{2}O)_{2} and (^{14}N^{15}NO)(^{15}N_{2}O) are assigned and fitted to Hamiltonians allowing rotational, centrifugal distortion and ^{14}N nuclear quadrupole coupling constants to be determined. Hyperfine structure is assigned for a single J′_{K′_{−1}K′_{+1}}→J′′_{K′′_{−1}K′′_{+1}} transition of both isotopomers of (^{14}N_{2}O)(^{15}N_{2}O). Nuclear quadrupole coupling

constants,χ_{bb}, are reported for all four ^{14}N nuclei. The measured χ_{bb} are in excellent agreement with those structures predicted from the measured rotational constants. The geometry of the molecule is slipped-parallel. The separation between the central nitrogen nuclei of the monomers in the r_{m}^{(1)} structure is 3.570(12)Å with the two N_{2}O monomers respectively oriented 54.69(68)^{∘} and 49.85(64)^{∘} to the a-inertial axis. Simulation of hyperfine structure in the spectrum of the (^{14}N_{2}O)_{2} isotopomer yields good qualitative agreement with experiment.